# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v3

@@ Abstract
The Lucas critique emphasizes the importance of considering microfoundations—how micro-agents (i.e., households) respond to policy changes—in macroeconomic policymaking. However, due to the vast scale and complex dynamics among micro-agents, predicting microfoundations is challenging. Consequently, this paper introduces a Stackelberg Mean Field Game (SMFG) approach that models macroeconomic policymaking based on microfoundations, with the government as the leader and micro-agents as dynamic followers. This approach treats large-scale micro-agents as a population, to optimize macroeconomic policies by learning the dynamic response of this micro-population. Our experimental results indicate that the SMFG approach outperforms real-world macroeconomic policies, existing AI-based and economic methods, enabling the learned macroeconomic policy to achieve the highest performance while guiding large-scale micro-agents toward maximal social welfare. Additionally, when extended to real-world scenarios, households that do not adopt the SMFG policy experience lower utility and wealth than adopters, thereby increasing the attractiveness of our policy. In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policymaking issues.
Figure 1: The mapping from macroeconomic policymaking (left) to Dynamic Stackelberg Mean Field Games (right). On the left, macro-level government actions, including fiscal or monetary policies, aim to optimize GDP, welfare, or equity. The microfoundations reflect the economic behaviors of large-scale micro-agents responding to shifts in macroeconomic policy. In Dynamic SMFGs, the government acts as the leader, and micro-agents as followers, respond to the leader’s policy.
@@ 1 Introduction
Macroeconomic policymaking is fundamental to the sustained development of an economy (Schneider and Frey 1988; Persson and Tabellini 1999). Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending (Sachs and Warner 1995). Therefore, accurately modeling and solving optimal macroeconomic policies, and evaluating the effects of their implementation are critical issues. According to the Lucas critique (Lucas Jr 1976), predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behaviors—termed “microfoundations”—to forecast micro-level responses to policy shifts and aggregate these into macroeconomic effects. Nonetheless, the vast number of micro-agents (i.e., households) and the complex dynamic interactions among them render predictions based on microfoundations exceedingly challenging.
To address the challenges outlined, this paper introduces a Stackelberg Mean Field Game (SMFG) approach: initially proposing a Dynamic Stackelberg Mean Field Game (Dynamic SMFG) framework to model macroeconomic policymaking based on microfoundations (see Figure 1). Within this framework, we design a model-free reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), aimed at optimizing macroeconomic policies by learning the dynamic responses of a large-scale population of micro-agents. In the Dynamic SMFG framework, the government acts as the leader, iteratively adjusting macroeconomic policies, while a vast array of micro-agents, serving as followers, react accordingly. This process entails sequential decision-making, with both the leader and the followers striving to optimize their cumulative long-term benefits. Given the leader’s policy, the SMFG is reduced to a mean field game, treating the large-scale micro-agents as a population to study their dynamic response to policy shifts. This dynamic SMFG framework responds to the Lucas critique. Existing models often simplify Dynamic SMFGs into single-step decisions (Pawlick and Zhu 2017), repeated games (Alcantara-Jimenez and Clempner 2020), or scenarios with myopic followers (Zhong et al. 2021), which fail to capture the complexities of macroeconomic policymaking adequately. Therefore, we introduce the SMFRL algorithm to optimize long-term utility for both the macro-agent and numerous micro-agents under the dynamic SMFG framework. In the SMFRL algorithm, we utilize mean field approximation to reduce the dimensionality of the micro-agents’ high-dimensional joint states and actions and employ a leader-follower update to learn the micro-population’s best response to macroeconomic policy adjustments, as well as to optimize macroeconomic policies considering dynamic feedback from the micro-population. This learning method enhances the stability and convergence of policies without requiring transition information or prior knowledge of the environment.
Our experiments, conducted on the TaxAI (Mi et al. 2023) environment, firstly demonstrate SMFG approach’s superiority over AI Economist (Zheng et al. ), the Saez tax (Saez 2001), the 2022 U.S. Federal tax, and free market policy (Backhouse 2005) from perspectives of macroeconomic policy evaluation and dynamic responses to economic shocks. Further, through ablation studies, we verify the significant roles of mean field approximation and leader-follower update in terms of macroeconomic policy effects, convergence, and maximal social welfare. When applied in real-world scenarios, the SMFG approach shows that households refusing the policy have lower wealth and utility than adopters, which enhances the SMFG policy’s attractiveness.
In summary, this paper’s contributions are threefold:
1. We introduce the Dynamic Stackelberg Mean Field Games framework, a novel approach for modeling macroeconomic policymaking based on microfoundations, responding to the Lucas critique.
2. We propose a model-free Stackelberg Mean Field Reinforcement Learning algorithm, designed to optimize macroeconomic policies by learning the dynamic responses of large-scale micro-agents within the Dynamic SMFG framework.
3. Experimental results demonstrate the superiority of our method in modeling and solving macroeconomic policymaking, and further explore its effectiveness in real-world settings.
Our code is available in an anonymous GitHub repository
https://anonymous.4open.science/r/SMFG˙macro˙7740
@@ 2 Related work
Economic Methods
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model (Hicks 1980; Gali 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model (Dutt 2006; Lee, Pesaran, and Smith 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model (Brock and Taylor 2010) provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models (Blanchard and Galí 2007; Gabaix 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models (An and Schorfheide 2007; Smets and Wouters 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis (Ramesh et al. 2010; Vedung 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods (Johnston and DiNardo 1963; Davidson, MacKinnon et al. 2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality. Additionally, the Saez tax model (Saez 2001) uses earnings elasticities to optimize tax rates, enhancing social welfare without deterring economic activity. This model continues to guide current fiscal policy through effective taxation strategies.
AI for Economics
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury 2022). In terms of macroeconomics, AI Economist (Zheng et al. ) employs curriculum learning on tax policy design, (Trott et al. 2021) studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang and Tänzer ; Chen et al. 2023), international trade (Sch 2021), market pricing (Danassis et al. 2023). However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. (Koster et al. ) studies democratic AI by human voting models, (Yaman et al. 2023) examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, (Shi 2021; Rui and Shi 2022; Atashbar and Aruhan Shi 2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha 2021; Hill, Bardoscia, and Turrell 2021), multiple equilibria under real-business-cycle (Curry et al. 2022), the emergence of barter behavior (Johanson et al. 2022), optimal asset allocation and savings strategies (Ozhamaratli and Barucca 2022).
Stackelberg Mean Field Game
Stackelberg Mean Field Games (SMFGs) model the dynamic interactions between a leader and numerous homogeneous followers, enabling the development of macroeconomic policies grounded in microfoundations. The methods for SMFGs are divided into model-based and model-free algorithms. Model-based methods, such as (Bensoussan, Chau, and Yam 2015; Fu and Horst 2020; Dayanikli and Lauriere 2023; Bergault, Cardaliaguet, and Rainer 2023) employ forward-backward stochastic functional differential equations to ascertain the followers’ mean field equilibrium prior to computing the optimal strategy for the leader. Additionally, studies such as (Bensoussan et al. 2017; Lin, Jiang, and Zhang 2018; Moon and Başar 2018; Du, Wu et al. 2019; Huang and Yang 2020) simplify scenarios through linear-quadratic formulations. (Guo, Hu, and Zhang 2022), convert SMFGs into minimax optimization problems. However, these model-based methods generally depend on simplified representations of SMFGs and explicit environmental transition information. In contrast, model-free approaches leverage interaction data with the environment to learn strategies without requiring detailed knowledge of the transition dynamics and reward functions.(Pawlick and Zhu 2017) explores the resolution of SMFGs in single-step decision scenarios.(Campbell et al. 2021) utilizes a modified deep backward stochastic differential equation framework to calculate the followers’ equilibrium despite constraints on the leader’s decisions.(Miao et al. 2024) investigates locally optimal defensive strategies given predefined attacker trajectories. Furthermore,(Li et al. 2024) learns followers’ transition functions from empirical data, thus aiding the solution of SMFGs through the empirical Fokker-Planck equation. Nevertheless, the assumptions and constraints inherent in these algorithms often do not meet the complex needs of macroeconomic policymaking.
3 Stackelberg Mean Field Game Modeling
In this section, we provide details about macroeconomic policy-making problems and propose a modeling framework based on Dynamic Stackelberg Mean Field Games (Dynamic SMFGs), with the relevant mappings are shown in Figure 1.
Macroeconomic Policy-Making
To foster economic growth, enhance social welfare, or maintain social equity, governments implement economic policies such as fiscal and monetary policies. These policies directly influence micro-agents (i.e., households) and the markets they constitute, altering their expectations and behaviors. The aggregation of micro-level behaviors from numerous households manifests as macroeconomic phenomena. Therefore, in formulating policies, governments should model micro-agents’ behaviors and predict their responses to new policies, which is so-called microfoundations in the Lucas critique. However, in reality, micro-agents form a large and dynamically interacting group, each seeking to maximize personal benefits. This complexity significantly complicates predicting their responses to macroeconomic policy shifts.
To address these challenges, we propose using Dynamic SMFGs to model macroeconomic policy-making problems:
1. For microfoundations, we model the government as the leader who first sets policies, with micro-agents as followers who adjust their behaviors according to these policies. The interaction between the leader and followers forms a Stackelberg game, where the leader considers the followers’ responses to make informed decisions.
2. For large scale micro-agents, we employ a Mean Field Game to model micro-interactions among a large number of households, treating them as a population and predicting their collective response to macroeconomic policy changes.
3. For dynamic decision-making, the formulation of macroeconomic policies is inherently a multi-step decision-making process, where both the government and households make decisions at each step to optimize long-term benefits, thereby constituting a dynamic game.
Dynamic Stackelberg Mean Field Games
@@ 3 Stackelberg Mean Field Game Modeling
For macroeconomic policy-making issues, we consider a dynamic SMFG with a leader and $N$ homogeneous follower agents. At any time step $t\in\{0,\ldots,T\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$. Subsequently, followers determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$. The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively. We denote the population state-action distribution of followers under leader action $a^{l}_{t}$ and follower policy $\pi^{f}$ as $L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})$, and abbreviate it as $L_{t}(s^{f}_{t},a^{f}_{t})$.
$L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f}),\text{where }a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$
Followers’ Mean Field Game
@@
Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers. At time $t\in\{0,\ldots,T-1\}$, given $(a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})$, the representative follower then receives a reward $r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$ and transitions to the next state $s^{f}_{t+1}$ according to the transition probability $s^{f}_{t+1}\sim P(\cdot|\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$, where joint state $\mathbf{s}_{t}=\{s^{l}_{t},\mathbf{s}_{t}^{f}\}$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})\right]$
@@
Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distributions $L=\{L_{t}\}_{t=0}^{T}$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L)$ is defined as
@@
Leader’s Game
At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, the followers’ policy $\pi^{f}$ and state-action distribution $L_{t}$, the leader will receive a reward $r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})$, and move to the next state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|\mathbf{s}_{t},a_{t}^{l},L_{t})$. The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
@@
In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic SMFG. Under this modeling framework, we can optimze the macroeconomic policy of the leader, namely the government, while considering the followers’ best response over discrete timesteps (Definition 2).

### v4

# Learning Macroeconomic Policies through Dynamic Stackelberg Mean-Field Games

@@ Abstract
Macroeconomic outcomes emerge from individuals’ decisions, making it essential to model how agents interact with macro policy via consumption, investment, and labor choices. We formulate this as a dynamic Stackelberg game: the government (leader) sets policies, and agents (followers) respond by optimizing their behavior over time. Unlike static models, this dynamic formulation captures temporal dependencies and strategic feedback critical to policy design. However, as the number of agents increases, explicitly simulating all agent–agent and agent–government interactions becomes computationally infeasible. To address this, we propose the Dynamic Stackelberg Mean Field Game (DSMFG) framework, which approximates these complex interactions via agent–population and government–population couplings. This approximation preserves individual-level feedback while ensuring scalability, enabling DSMFG to jointly model three core features of real-world policymaking: dynamic feedback, asymmetry, and large scale. We further introduce Stackelberg Mean Field Reinforcement Learning (SMFRL), a data-driven algorithm that learns the leader’s optimal policies while maintaining personalized responses for individual agents. Empirically, we validate our approach in a large-scale simulated economy, where it scales to 1,000 agents (vs. 100 in prior work) and achieves a 4×4$\times$4 × GDP gain over classical economic methods and a 19×19$\times$19 × improvement over the static 2022 U.S. federal income tax policy.
[…]
3449
@@ 1 Introduction
Macroeconomic policy formulation is fundamental to achieving sustainable economic growth [52, 46]. The effectiveness of these policies depends crucially on the behaviors of micro-level individuals, such as labor supply, consumption, and investment decisions [49]. Nobel laureate Lucas has emphasized that individuals adapt their decision-making in response to changes in macroeconomic policies [37]. Thus, systematically modeling the interactions between individuals and the government is crucial for designing effective macroeconomic policies.
The Stackelberg game naturally captures the asymmetric interactions [19, 30] where the government (leader) sets a tax policy, and individuals (followers) adjust their labor supply and consumption in response. Unlike static models, this dynamic formulation accounts for the time-dependent strategic feedback essential for effective policy design [43, 5, 15]. However, scaling this approach to large populations is computationally intractable: with $N$ agents, there are $O(N^{2})$ pairwise agent-agent interactions plus $O(N)$ leader–agent interactions.
To address the scalability challenge inherent in macroeconomic policy modeling, we propose the Dynamic Stackelberg Mean Field Game (DSMFG) framework, which unifies dynamic feedback, asymmetry, and large scale into a coherent formulation. Unlike standard SMFG methods—e.g., single-step models [19] or approaches with fixed dynamics [8]—DSMFG captures the multi-period feedback loops essential to the co-evolution of policy and individual response. By embedding a multi-step Stackelberg game within a mean-field approximation [12, 1], DSMFG reduces the $O(N^{2})$ complexity of agent-agent interactions to $O(N)$ agent–population interactions. At each timestep, the government optimizes policy based on the current mean field—i.e., the population’s state–action distribution—while agents adapt to both the policy and the mean field. This iterative structure preserves individual feedback, reduces computational cost, and enables scalable optimization in complex macroeconomic environments.
To learn optimal policies under DSMFG, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm. SMFRL introduces a Stackelberg Mean Field Q-function that enables the leader to evaluate its interactions with the aggregate population. Another Q-function for followers evaluates their interactions with both the leader and the population. Followers share a common policy that takes heterogeneous features as input, enabling personalized actions while maintaining population-level consistency. The central Q and shared policy are designed to ensure scalability in large populations. Moreover, SMFRL employs alternating updates between the leader’s and followers’ policies to ensure a stable training.
We empirically evaluate DSMFG and SMFRL in a large-scale macroeconomic simulation environment, TaxAI, which models dynamic interactions between the government and large scale agents. Comparing against static policies (e.g., the 2022 U.S. federal tax) and dynamic rule-based methods (e.g., the Saez tax), DSMFG yields substantially better outcomes—achieving a 4×4$\times$4 × gain in per capita GDP over the Saez tax and a 19×19$\times$19 × improvement over the 2022 U.S. baseline. Unlike static or independent-agent baselines, DSMFG ensures sustainability and rapidly stabilizes income and consumption after shocks. Ablation studies confirm the necessity of both Stackelberg hierarchy and mean-field approximation: removing either leads to lower welfare, instability, and poor convergence. Despite sharing a policy, followers maintain robust performance under population heterogeneity, enabling scalable training and consistently high utility. These results demonstrate DSMFG’s effectiveness in solving dynamic, large-scale government–agent problems within a controlled, reproducible simulation framework, extending beyond the scope of traditional approaches.
In summary, Our key contributions are:
Figure 1: Macroeconomic policy-making involves both intensive individual–individual and individual–government interactions, whose pairwise complexity grows as $O(N^{2})$ with the agent number $N$—rendering direct simulation intractable. We address this by modeling the process as a Dynamic Stackelberg Mean Field Game (DSMFG), which approximates these complex interactions via agent–population and leader–population couplings. This DSMFG retains key personal behaviors while enabling scalability and capturing three defining features of real-world policymaking: dynamics, asymmetry, and large-scale.
  • A scalable DSMFG framework that integrates three core features—dynamic feedback, asymmetry, and large scale—into a unified model for macroeconomic policymaking. (Section 3)
  • A SMFRL algorithm that efficiently learns macro policies under DSMFG while preserving personalized decision-making at the individual level. (Section 4)
  • Extensive empirical validation showing that SMFG scales to 1,000 agents and outperforms both economic and AI-based baselines. (Section 5)
@@ 2 Related Work
2.1 Macroeconomic Models
Classical macroeconomic frameworks—such as IS–LM [26, 22], AD–AS [20, 34], and the Solow growth model [10]—have elucidated short- and long-run policy effects. New Keynesian DSGE models [9, 56] then introduced micro individuals and stochastic shocks, enabling rigorous analysis under rational expectations. The Saez tax model [50] further offered a practical, elasticity-based tool for setting optimal tax rates. However, these approaches commonly rely on linearization, representative-agent assumptions, or fixed price-stickiness parameters, which prevent them from capturing individual responses, nonlinear feedback loops, and aggregate dynamics in large populations. Empirical and econometric methods [47, 17] quantify policy impacts from historical data but struggle with sparsity, identification, and out-of-sample validity. These gaps highlight the necessity of a framework that models dynamic interactions between the government and large-scale individuals.
2.2 RL for Economic Policy
Recent advances have applied reinforcement learning (RL) into economic modeling. In macroeconomic settings, frameworks like AI Economist [60] use curriculum learning to optimize tax schedules, and others apply RL to crisis management [57], monetary policy design [28, 14], international trade dynamics [51], and market pricing with externalities [16]. These studies typically treat the government as an independent decision-maker, overlooking how heterogeneous households adapt over time. At the micro level, RL has been used to study optimal savings and consumption [53, 48, 3], solve heterogeneous general-equilibrium problems [32, 27], and model agent behaviors in barter [31] and asset allocation [44]. While these works showcase RL’s promise, they rely on simplified environments or small agent populations, limiting their applicability to dynamic, large-scale macroeconomic policy design.
2.3 Stackelberg Mean Field Games
Stackelberg Mean Field Games (SMFGs) integrate Stackelberg leader–follower dynamics with mean-field approximations to model interactions in large populations. Early model-based SMFG methods solve forward–backward stochastic differential equations to compute follower equilibria before optimizing the leader’s policy [21, 18, 8]. Linear–quadratic formulations [7, 42, 29] and minimax rewritings [23] enhance analytical tractability but impose restrictive assumptions on dynamics and transitions, limiting applicability to complex economic settings. Model-free SMFG approaches dispense with explicit transition models by learning directly from interaction data. For example, Pawlick and Zhu [45] handle single-step SMFGs, Campbell et al. [11] apply deep BSDE solvers for equilibrium computation, and Miao et al. [40] explore defensive follower strategies under fixed attacker trajectories. More recently, Li et al. [35] estimate empirical transition kernels and solve the resulting Fokker–Planck equations. However, these methods remain confined to simplified benchmarks—single-period decisions, low-dimensional state spaces, or predefined follower classes—and do not capture the multi-period feedback loops and scale required for realistic macroeconomic policy design. Thus, there is a clear need for a stable, model-free algorithm that can solve SMFGs under complex, dynamic economic interactions without requiring knowledge of true transition dynamics.
3 Dynamic Stackelberg Mean Field Game Framework
In this section, we first identify the core features of the macroeconomic policy-making problem, and then propose a dynamic Stackelberg mean-field game framework to model them effectively.
3.1 Macroeconomic Policy-Making Problem
Macroeconomic policy comprises government actions—such as monetary and fiscal interventions—that stabilize growth, reduce unemployment, and control inflation [6]. These interventions shape individual decisions (e.g. labor supply, consumption, investment), which in turn generate aggregate outcomes that inform subsequent policy adjustments [41]. In the left panel of Figure 1, for instance, a change in the central bank’s interest-rate rule shifts households’ portfolio allocations, while each household’s choice also depends on the behavior of others. Such large-scale interactions induce complex feedback loops that static or small-scale models fail to capture.
We identify three salient, interdependent features of the macroeconomic policy-making problem:
  1. 1.
    Dynamic feedback. A policy change triggers micro-level behavioral adjustments; the aggregate of these adjustments produces new macro indicators, which then feed back into the next policy decision. Modeling this continuous loop is essential, yet beyond the scope of static modeling methods.
  2. 2.
    Asymmetry. The government (leader) first commits to a policy rule; individual agents (followers) observe this policy and then optimize their private objectives. This sequential leader-follower structure underlies the inherent asymmetry dynamics between policymaker and population.
  3. 3.
    Large scale. Effective macro policy influences large scale micro-agents, thereby rendering the dynamic, asymmetric interactions described above more complex.
3.2 Dynamic Stackelberg Mean Field Game
To design effective macroeconomic policies, we must first model three core features—dynamic feedback, asymmetry, and large scale—introduced above.
To capture dynamic feedback, we model macroeconomic policymaking as a dynamic game in which the government and individuals iteratively update their strategies based on evolving economic conditions. These strategies influence both immediate outcomes and the future decisions of other players [25, 58].
To capture asymmetry, we adopt a Stackelberg leader–follower framework [55], modeling the government as the leader that sets policies first, followed by individuals’ responses.
To capture the large-scale nature of macroeconomic systems, an agent-based model with numerous micro-agents could be considered. However, scaling agent-based models to large populations is computationally challenging. With $N$ agents, the model requires $O(N^{2})$ pairwise agent–agent interactions and $O(N)$ government–agent interactions, rendering it computationally intractable for large-scale systems. To address this challenge, we employ a mean-field approximation [33], where interactions between individual agents are replaced by those between a representative agent and the aggregate population, and government–agent interactions are modeled as interactions with the population’s mean field. This approach reduces the computational complexity from $O(N^{2})$ to $O(N)$, significantly streamlining policy optimization.
In summary, our Dynamic Stackelberg Mean-Field Game (DSMFG) framework effectively integrates the three core features—dynamic feedback, asymmetry, and large scale—into a cohesive model for macroeconomic policymaking. The precise mathematical formulation of DSMFG is as follows:
Framework Overview
@@ 3.2 Dynamic Stackelberg Mean Field Game
In the DSMFG framework, we consider one leader and $N$ follower agents. At each time step $t\in\{0,\ldots,T\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on its state $s_{t}^{l}\in\mathcal{S}^{l}$ and a policy $\pi^{l}:\mathcal{S}^{l}\to\mathcal{A}^{l}$. Subsequently, the followers determine their actions based on the leader’s action $a_{t}^{l}$ and their private states $s_{t}^{f}\in\mathcal{S}^{f}$. A representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\to\mathcal{A}^{f}$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively. The mean field $L_{t}(s_{t}^{f},a_{t}^{f};\pi^{f},a_{t}^{l})$, abbreviated as $L_{t}(s_{t}^{f},a_{t}^{f})$, represents the population state-action distribution of followers, defined as:
$L_{t}(s_{t}^{f},a_{t}^{f};\pi^{f},a_{t}^{l})\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f}),\text{ where }a_{t}^{f}=\pi^{f}(s_{t}^{f},a_{t}^{l}).$
Followers
@@
At each time step $t\in\{0,\ldots,T-1\}$, given the joint state $\mathbf{s}_{t}=\{s_{t}^{l},\mathbf{s}_{t}^{f}\}$, a representative follower receives a reward $r^{f}(\mathbf{s}_{t},a_{t}^{l},a_{t}^{f},L_{t})$ and transitions to the next state $s_{t+1}^{f}\sim P(\cdot\mid\mathbf{s}_{t},a_{t}^{l},a_{t}^{f},L_{t})$. The follower’s objective is to optimize their policy $\pi^{f}$ to maximize cumulative rewards over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}r^{f}(\mathbf{s}_{t},a_{t}^{l},a_{t}^{f},L_{t})\right],$
where the leader’s action $a_{t}^{l}=\pi^{l}(s_{t}^{l})$, the follower’s action $a_{t}^{f}=\pi^{f}(s_{t}^{f},a_{t}^{l})$, and the mean field $L_{t}=L_{t}(s_{t}^{f},a_{t}^{f})$.
@@
Given a leader’s policy $\pi^{l}\in\Pi^{l}$ and followers’ state-action distributions $L=\{L_{t}\}_{t=0}^{T}$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L)$ is defined as:
@@
Leader
At each time step $t\in\{0,\ldots,T-1\}$, the leader receives a reward $r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})$ based on its state $s_{t}^{l}$, action $a_{t}^{l}$, and the population mean field $L_{t}$, and transitions to the next state $s_{t+1}^{l}\sim P(\cdot\mid\mathbf{s}_{t},a_{t}^{l},L_{t})$. The leader aims to optimize its policy $\pi^{l}$ to maximize the expected cumulative reward:
@@
In conclusion, we model the problem of macroeconomic policy-making as a DSMFG, capturing three core features. Based on this model, we optimize the government’s policy by taking into account the followers’ best responses over discrete timesteps (Definition 2).